OpenStudy (anonymous):

Find the point on the terminal side of θ = negative three pi divided by four that has an x coordinate of -1

OpenStudy (jdoe0001):

 recall your SOH CAH TOA thus \(\bf tan\left(-\frac{3\pi}{4}\right)=\cfrac{opposite}{adjacent}\to \cfrac{y}{x}\implies x\cdot tan\left(-\frac{3\pi}{4}\right)=y\)

OpenStudy (jdoe0001):

or well since x = -1 then \(\bf -1\cdot tan\left(-\frac{3\pi}{4}\right)=y\)